OpenStudy (anonymous):

Let f(x) = x^3 + bx, where b is an element of the real numbers. Find all value(s) of b such that f`(x)=0 for all real values of x.

OpenStudy (bahrom7893):

f'(x) = 0 means: f'(x) = 3x^2+b = 0

OpenStudy (bahrom7893):

3x^2 + b = 0 3x^2 = -b So now, since the left side can only be a positive number.

OpenStudy (bahrom7893):

And we're only dealing with reals in here, b can range from -infinity to 0, or: (-inf, 0]

OpenStudy (bahrom7893):

basically b has to be non-positive.

OpenStudy (anonymous):

hm..i don't see why

OpenStudy (bahrom7893):

because 3*x^2 cannot equal to a negative number

OpenStudy (bahrom7893):

3 = positive x^2 = positive (any number squared is positive) positive*positive = positive 3*(x^2) = -b

OpenStudy (bahrom7893):

thus -b must be positive, or b must be negative.. here let me use an example.

OpenStudy (bahrom7893):

Suppose b = 1 3x^2 = -1 x^2 = -1/3 but we are only dealing with reals, not imaginary numbers. That's why b cannot be positive.

OpenStudy (bahrom7893):

Now suppose b = 0 3*x^2 = 0 x^2 = 0, x = 0. So 0 works.

OpenStudy (bahrom7893):

Now suppose b = -1 3*x^2 = -(-1) 3x^2 = 1 x^2 = 1/3 x = +/- sqrt(1/3) See. This only works with 0 or negative numbers.

OpenStudy (anonymous):

okay. i am following what you're saying. just have to form it into words. my professor is really strict with forming our proofs.

OpenStudy (bahrom7893):

WAIT a second lol

OpenStudy (anonymous):

what'd i do? haha

OpenStudy (bahrom7893):

i forgot, it didn't just say exists, it said = 0.

OpenStudy (anonymous):

ok...

OpenStudy (bahrom7893):

oh nevermind i was right, i did take that into account.

OpenStudy (anonymous):

did it matter that you solved for b though when the question asked for all real values of x? is it referring to once you've solved the equation completely?

OpenStudy (bahrom7893):

x is real means that that it's square cannot be negative.

OpenStudy (bahrom7893):

that's why I knew that b had to be negative.

OpenStudy (bahrom7893):

Okay let's start over: f'(x) = 3x^2 + b = 0 3x^2 + b = 0 3x^2 = -b x^2 = -b/3 So far so good, right?

OpenStudy (bahrom7893):

x is a real number, therefore \[x^2 \ge 0\] So: \[-b/3 \ge 0\]

OpenStudy (anonymous):

okay...now it's in terms of x.

OpenStudy (bahrom7893):

Multiplying both sides of the equation by 3 gives: \[-b \ge 0\] Now multiply both sides by -1, but remember when multiplying by a negative sign, u reverse the inequality, so you end up with: \[b \le 0\]

OpenStudy (bahrom7893):

good?

OpenStudy (anonymous):

good. back to my last question...i wrote a proof out. would you let me know if it makes any sense

OpenStudy (bahrom7893):

kk hold on